A peregrine falcon is the fastest animal in the world. When it dives, it can accelerate to about 40m/s. Calculate the length of the time the bird needs to get to its top speed from rest by taking it's acceleration to be 8 m/


1
Expert's answer
2021-08-26T10:24:47-0400

"v=v_0+at=0+8\cdot t\to40=8\cdot t\to t=5\ (s)" . Answer